What should be tension applied to a wire of length 1m and mass 10g if it has to vibrate with fundamental frequency of 50 Have?

Grade:12

1 Answers

Arun
6 years ago
 
Dear Akshay
 
mass m = 10 gm = 0.01
n = (1/2L) \sqrt(T/m)
Now
50 = 1/2 \sqrt(T/0.01)
100² = T/0.01
10000*0.01 = T
T = 100 N
 
Regards
Arun (askIITians forum expert)